There are three (3) cars moving differently with respect to the earth. Suppose that Car A is moving 20km/h

20km/h to the east, Car B is moving 8km/h

8km/h to the east, while Car C is moving 5km/h

5km/h to the west.

a. What is the velocity of Car A with respect to Car B?

b. What is the velocity of Car C with respect to Car B?



1
Expert's answer
2021-12-02T10:04:43-0500

a. Car A and car B are moving in the same direction, so, their relative velocity (of A with respect to B) is


"v_{AB}=(20-8)=12\text{ km/h East}."

b. The velocity of Car C with respect to Car B:


"v_{CB}=(5+8)=13\text{ km/h West.}"